OpenStudy (anonymous):

Please help!! Simplify the expression. (6 - square root 11) (6 + square root 11)

OpenStudy (anonymous):

\[(6 -\sqrt{11}) ( 6 + \sqrt{11}\]

OpenStudy (richyw):

\[(6-\sqrt{11})(6+\sqrt{11})\]\[36-6\sqrt{11}+6\sqrt{11}-11\]\[36-11=25\]

OpenStudy (campbell_st):

this is the difference of 2 squares the solution when expanded is \[6^2 - (\sqrt{11})^2 \] you just need to evaluate expression for the answer

OpenStudy (anonymous):

Thank you both very much!